# Boil-off of red supergiants: mass loss and type II-P supernovae

### v1

@@ Abstract
The mass loss mechanism of red supergiant stars is not well understood, even though it has crucial consequences for their stellar evolution and the appearance of supernovae that occur upon core-collapse. We argue that outgoing shock waves launched near the photosphere can support a dense chromosphere between the star’s surface and the dust formation radius at several stellar radii. We derive analytic expressions for the time-averaged density profile of the chromosphere, and we use these to estimate mass loss rates due to winds launched by radiation pressure at the dust formation radius. These mass loss rates are similar to recent observations, possibly explaining the upward kink in mass loss rates of luminous red supergiants. Our models predict that low-mass red supergiants lose less mass than commonly assumed, while high-mass red supergiants lose more. The chromospheric mass of our models is $\sim$0.01 solar masses, most of which lies within a few stellar radii. This may explain the early light curves and spectra of type-II P supernovae without requiring extreme pre-supernova mass loss. We discuss implications for stellar evolution, type II-P supernovae, SN 2023ixf, and Betelgeuse.
@@ 1 Introduction
It is widely accepted that these winds are accelerated by radiation pressure on dust (e.g., Hoyle & Wickramasinghe 1962), which forms well above their photospheres where the temperature drops to $T\lesssim 1500\,{\rm K}$ (Field 1974; Höfner & Olofsson 2018) such that dust can form (Figure 1). The high opacity of dust means that dust-laden gas has a large enough opacity to exceed the Eddington limit, such that it will be accelerated outwards by the momentum of photons emitted from the star. However, the dust can only form at radii several times larger than the photosphere of the star, $R_{\rm d}\gtrsim 5\,R$, and the mechanisms by which material is lofted up to the dust formation radius are highly uncertain. This wind-launching problem for RSGs is crucial for determining their mass loss rates, and it is the main focus of this paper.
Perhaps the most detailed investigations of mass loss from cool stars have been performed for asymptotic giant branch (AGB) stars (see Höfner & Olofsson 2018 for a review). Although these stars have lower mass and smaller luminosities than RSGs, they have similar surface temperatures and escape speeds, and the mass loss mechanism is believed to be similar to that described above. For AGB stars, some combination of turbulent pressure and shock waves driven by radial pulsations are thought to propel gas above the photosphere (Höfner et al., 2003). Because this chromospheric material is dense, it can cool (unlike the Sun’s corona, Cranmer & Saar 2011) so that its temperature maintains radiative equilibrium with the star and decreases approximately as $T\propto r^{-1/2}$ in the chromosphere. Evermore sophisticated radiation hydrodynamics simulations of this process (Freytag & Höfner, 2008; Freytag et al., 2017; Freytag & Höfner, 2023) support this general picture and also give detailed predictions for atomic, molecular, and dust spectral features in the spectra of these stars.
@@
This CSM is orders of magnitude denser than expected from stellar evolution models of RSGs, which predict sharp drops in density above the photosphere. The CSM is also a few orders of magnitude denser than expected if the CSM is formed by a wind with mass loss rate and velocity typical for RSGs, $\dot{M}\sim 10^{-6}\,M_{\odot}/{\rm yr}$, $v_{\rm wind}\sim 30\,{\rm km}/{\rm s}$, (Dessart et al., 2017; Moriya et al., 2017; Morozova et al., 2017; Moriya et al., 2018; Morozova et al., 2018). The frequent conclusion of SN observers and modelers has thus been that RSGs lose mass at much higher rates during the final years of their lives. This idea has been supported by some RSG models that have predicted pre-supernova outbursts that eject mass in the weeks-months before explosion (Woosley & Heger, 2015; Fuller, 2017).
@@ 2.1 Equations of Motion
The $\rho{\vec{v}}$ term is proportional to the mass flux, which may be non-zero but is assumed to be constant in space/time. Hence, averaging over time, this term vanishes. For simplicity, we will neglect thermal pressure since it is small for moderate or strong shocks (see Bertschinger & Chevalier 1985) for a model including thermal pressure). We also neglect radiative forces because the low opacity and density of the chromosphere causes it to be optically thin, but later we will account for radiative acceleration of dust grains in the dust formation region.
@@
Another justification for a nearly constant shock speed above the photosphere is the isothermal nature of the shocks produced. Following Matzner & McKee (1999), a shock wave propagating through a star accelerates the material to a velocity of
@@
A third justification for a constant shock speed as a function of radius are the models of Bertschinger & Chevalier (1985), who examined atmospheres of Mira variables supported by a periodic train of outgoing shocks. In the plane-parallel limit, they showed that periodic solutions have shock velocities that are constant as a function of height, while the density falls exponentially (when averaging over time), as shown by the calculation below. The plane-parallel limit applies when the separation between subsequent shocks is less than the stellar radius, which is only marginally true in our case. Simulations (e.g., Goldberg et al. 2022a) show that shocks are launched approximately on the convective turnover time $t_{\rm con}\sim 0.5R/v_{\rm con}$ such that the separation between shocks is $\sim 0.5R$.
@@ 2.2 Chromospheric Structure
This is similar to the structure of an isothermal atmosphere, but with the sound speed replaced by $v_{\rm s}$, and with the extra geometric factor of $(R/r)^{2}$ due to the use of the divergence of the momentum flux rather than a turbulent pressure. Just above the stellar surface, equation 5 is identical to the density profile found by Bertschinger & Chevalier (1985) for a periodic train of shocks in a plane-parallel atmosphere. Importantly, the density profile depends only on the amplitude of shocks, and not on their temporal frequency.
@@
Figure 2 shows the density profiles of equation 5 and equation 2.2 above a RSG photosphere. The underlying model is a $13.8\,M_{\odot}$ RSG with $R=844R_{\odot}$ generated using MESA (Paxton et al., 2011; Paxton et al., 2013, 2015). This model had a ZAMS mass of $15\,M_{\odot}$ and has been evolved to core-collapse. We use MLT with $\alpha_{\rm MLT}=2.5$ to measure the convective velocity, which we evaluate where the optical depth is $\tau=P_{\rm rad}c/(Pv_{\rm con})$, where $P_{\rm rad}$ and $P$ are the radiation and total pressures. This yields approximately correct RMS convective velocities (see Goldberg et al. 2022a), which for RSGs are quite similar to the sound speed $c_{s}$. This model has $v_{\rm con}\simeq 0.08v_{\rm esc}$ near the star’s surface.
@@ 3.1 Dust Formation and Wind Acceleration
We choose to use observed wind velocities $v_{\infty}$ to constrain $x_{\rm d}$, but one could use dust opacities and mass fractions to calculate $x_{\rm d}$ and then predict the corresponding value of $v_{\infty}$. From equation 14, we have $x_{\rm d}=(R_{\rm d}/GM)(v_{\infty}^{2}-v_{\rm d}^{2})$. In this work, we use $v_{\infty}=30$ km/s, in line with observational estimates (Sjouwerman et al., 1998; van Loon et al., 2001; Marshall et al., 2004), although this velocity likely depends weakly on the stellar luminosity (e.g., Mauron & Josselin 2011). This requires $x_{\rm d}\sim 2$, approximately what we expect for realistic dust-laden gas opacities of $\kappa_{\rm d}\sim 5\,{\rm cm^{2}}/{\rm g}$ (Höfner et al., 2003).
@@ 3.2 Red Supergiant Mass Loss Rates
We implement the mass loss rate of the previous section (equation 3.1) into stellar evolution models generated with MESA as described above. At surface temperatures above $4500\,{\rm K}$ we use the “Dutch” wind prescription with scaling factor $\eta=0.5$. At cooler surface temperatures, we use our wind prescription.
@@
Our mass loss rates are similar for a given luminosity to those recently measured for RSGs in the LMC (Antoniadis et al., 2024; Wen et al., 2024). These mass loss rates are substantially smaller than those measured in the SMC (Yang et al., 2023), but those may be systematically overestimated according to Antoniadis et al. 2024. Most importantly, the predicted scaling of mass loss with luminosity above $L>3\times 10^{4}\,L_{\odot}$ is quite similar between our models and all of these measurments. Interestingly, the LMC/SMC measurements indicate a kink at $L\sim 3\times 10^{4}\,L_{\odot}$, with the mass loss rate flattening at smaller luminosities, which does not happen in our models. The observed mass loss rates at low luminosities are closer to the predictions of a Reimer’s mass loss prescription. It is likely that a different mass loss mechanism (e.g., mass loss driven by coronal heating due to Alfvén waves, Cranmer & Saar 2011) dominates at low luminosities.
@@
We do not plot the mass loss prescription of Schröder & Cuntz (2005), which lies between the Reimers and de Jager prescriptions for our models.
@@ 3.3 Implications of mass loss
Our massive models can exhibit exponentially increasing mass loss rate with time. As these stars lose mass, their luminosities/radii remain nearly constant but their masses decrease, decreasing $v_{\rm esc}$ and increasing the mass loss rate, creating a runaway process. Combined with the rarity of such stars due to the IMF, this implies a correspondingly small number of observed very luminous RSGs with large mass loss rates, perhaps consistent with their low observed numbers (Beasor & Smith, 2022). In contrast, the Reimers prescription predicts much lower mass loss rates, longer RSG durations, and a larger number of observed very luminous RSGs. The boil-off phenomenon of our models can thus help explain the absence of RSGs observed with luminosities larger than $L_{\rm max}\approx 10^{5.5}\,L_{\odot}$ (Davies et al., 2018), while simultaneously preserving low-mass loss rates of less massive RSGs.
Vink & Sabhahit (2023) suggested that the upward kink in mass loss rates at high luminosity ($L\gtrsim 10^{4.5}\,L_{\odot}$) is caused by a transition from an optically thin to optically thick wind. This is motivated by an analogy with line-driven winds from hot stars, which show a similar feature as photons begin to scatter multiple times in optically thick winds, allowing them to drive higher mass loss rates. However, as far as we are aware, there is no model of radiatively driven winds in cool stars, e.g., due to photon pressure on molecular absorption lines. While photon pressure on dust helps accelerate RSG winds, the mass loss rate may instead be determined by how much mass can be uplifted to the dust formation radius, so it is not clear that optically thick dust absorption would actually drive higher mass loss rates. Our model can explain the dependence of mass loss on luminosity above the kink, but it cannot explain the dependence below the kink. We speculate that a different mechanism with weaker luminosity dependence dominates mass loss at luminosities below $10^{4.5}\,L_{\odot}$. Future work will be necessary to distinguish between our suggestion and that of Vink & Sabhahit (2023).
@@ 4 Effect on Type II-P Supernovae
Our model predicts typical densities of $\rho\gtrsim 10^{-14}\,{\rm g}/{\rm cm}^{3}$ within three stellar radii of RSGs. This is a few orders of magnitude larger than expected for constant velocity winds with similar mass loss rates (Figure 2). Therefore, we expect signs of CSM interaction in type-II SNe, even for “ordinary” pre-SN mass loss rates of $\sim\!10^{-6}-10^{-5}\,M_{\odot}/{\rm yr}$. As mentioned above, the total CSM mass of our models is in the range of $10^{-3}-10^{-1}\,M_{\odot}$, slightly smaller than estimates based on light-curve modeling of type-IIP SNe exhibiting signs of CSM interaction (Morozova et al., 2017; Dessart et al., 2017; Moriya et al., 2018). Our chromospheres qualitatively resemble the less massive but more extended CSM structure needed to match both the light curve and narrow emission lines of type II-P (Dessart & Jacobson-Galán, 2023), but detailed spectral modeling will be needed to determine whether they can reproduce observations. Here we model the impact of the chromosphere on the light curves of type II-P SNe.
@@
We do not perform spectral modeling here, but we note that non-LTE modeling of flash-ionized emission lines has constrained outflow velocities to low values of $v_{\rm wind}\lesssim 30\,{\rm km/s}$ in the case of SN 2013cu (e.g., Gräfener & Vink 2016). That event was a hydrogen-poor type-IIb SN for which our models are not directly applicable. In any case, our quasi-hydrostatic chromosphere models predict low outflow/shock speeds within the dust formation radius of exploding type II-P SN progenitors, corresponding to narrow widths of $v\lesssim 30\,{\rm km/s}$ of flash-ionized emission lines, although the widths may be increased by acceleration due to radiation pressure from the subsequent SN (Tsuna et al., 2023a).
@@ 4.1 Light curve modeling
When the SN shock propagates an extended dense CSM (e.g., the case for $v_{\rm s}/v_{\rm esc}=0.24$), the shock downstream is compressed into a thin dense shell due to radiative cooling. By default, SNEC obtains the “observed luminosity” by measuring the local luminosity at the photospheric radius where $\tau=2/3$. However, due to the thin-shell nature of the shocked region, we find numerical instabilities in the luminosity after the photosphere enters the shocked region (e.g., Fig 4 of Tsuna et al., 2023b). To avoid this, we instead measure the bolometric luminosity at the outermost part of the computational region ($\approx 3\times 10^{16}$ cm) where photons are free-streaming ($\tau\ll 2/3$). We use this luminosity and the photospheric radius, and adopt the same bolometric corrections as done in SNEC to obtain the V-band light curves.
@@
We note that the model V-band light curves can be affected by SNEC’s assumption of gas-radiation equilibrium and a fully thermalized radiation field. We may expect departures from thermal spectra for the luminosity generated around shock breakout, and that powered by prolonged CSM interaction (e.g., Nakar & Sari, 2010; Svirski et al., 2012; Morozova et al., 2018; Haynie & Piro, 2021; Tsuna et al., 2021). On the other hand, for the shock cooling phase after breakout the thermalization is expected to be more complete (Morozova et al., 2018). Additionally, the morphology of the light curve around the peak is possibly affected by the relatively low resolution within the breakout layer ($\tau\approx$ several tens), most notably for the ones with densest CSM.

### v2

@@ Abstract
The mass loss mechanism of red supergiant stars is not well understood, even though it has crucial consequences for their stellar evolution and the appearance of supernovae that occur upon core-collapse. We argue that outgoing shock waves launched near the photosphere can support a dense chromosphere between the star’s surface and the dust formation radius at several stellar radii. We derive analytic expressions for the time-averaged density profile of the chromosphere, and we use these to estimate mass loss rates due to winds launched by radiation pressure at the dust formation radius. These mass loss rates are similar to recent observations, possibly explaining the upward kink in mass loss rates of luminous red supergiants. Our models predict that low-mass red supergiants lose less mass than commonly assumed, while high-mass red supergiants lose more. The chromospheric mass of our models is $\sim$0.01 solar masses, most of which lies within a few stellar radii. This can help explain the early light curves and spectra of type-II P supernovae without requiring extreme pre-supernova mass loss. We discuss implications for stellar evolution, type II-P supernovae, SN 2023ixf, and Betelgeuse.
@@ 1 Introduction
It is widely accepted that these winds are accelerated by radiation pressure on dust (e.g., Hoyle and Wickramasinghe 1962), which forms well above their photospheres where the temperature drops to $T\lesssim 1500\,{\rm K}$ (Field 1974; Höfner and Olofsson 2018) such that dust can form (Figure 1). The high opacity of dust means that dust-laden gas has a large enough opacity to exceed the Eddington limit, such that it will be accelerated outwards by the momentum of photons emitted from the star. However, the dust can only form at radii several times larger than the photosphere of the star, $R_{\rm d}\gtrsim 5\,R$, and the mechanisms by which material is lofted up to the dust formation radius are highly uncertain. This wind-launching problem for RSGs is crucial for determining their mass loss rates, and it is the main focus of this paper.
Perhaps the most detailed investigations of mass loss from cool stars have been performed for asymptotic giant branch (AGB) stars (see Höfner and Olofsson 2018 for a review). Although these stars have lower mass and smaller luminosities than RSGs, they have similar surface temperatures and escape speeds, and the mass loss mechanism is believed to be similar to that described above. For AGB stars, some combination of turbulent pressure and shock waves driven by radial pulsations are thought to propel gas above the photosphere (Höfner et al., 2003). Because this chromospheric material is dense, it can cool (unlike the Sun’s corona, Cranmer and Saar 2011) so that its temperature maintains radiative equilibrium with the star and decreases approximately as $T\propto r^{-1/2}$ in the chromosphere. Evermore sophisticated radiation hydrodynamics simulations of this process (Freytag and Höfner, 2008; Freytag et al., 2017; Freytag and Höfner, 2023) support this general picture and also give detailed predictions for atomic, molecular, and dust spectral features in the spectra of these stars.
@@
This CSM is orders of magnitude denser than expected from stellar evolution models of RSGs, which predict sharp drops in density above the photosphere. The CSM is also a few orders of magnitude denser than expected if the CSM is formed by a wind with mass loss rate and velocity typical for RSGs, $\dot{M}\sim 10^{-6}\,M_{\odot}/{\rm yr}$, $v_{\rm wind}\sim 30\,{\rm km}/{\rm s}$, (Dessart et al., 2017; Moriya et al., 2017; Morozova et al., 2017; Moriya et al., 2018; Morozova et al., 2018). The frequent conclusion of SN observers and modelers has thus been that RSGs lose mass at much higher rates during the final years of their lives. This idea has been supported by some RSG models that have predicted pre-supernova outbursts that eject mass in the weeks-months before explosion (Woosley and Heger, 2015; Fuller, 2017).
@@ 2.1 Equations of Motion
The $\rho{\vec{v}}$ term is proportional to the mass flux, which may be non-zero but is assumed to be constant in space/time. Hence, averaging over time, this term vanishes. For simplicity, we will neglect thermal pressure since it is small for moderate or strong shocks (see Bertschinger and Chevalier 1985) for a model including thermal pressure). We also neglect radiative forces because the low opacity and density of the chromosphere causes it to be optically thin, but later we will account for radiative acceleration of dust grains in the dust formation region.
@@
Another justification for a nearly constant shock speed above the photosphere is the isothermal nature of the shocks produced. Following Matzner and McKee (1999), a shock wave propagating through a star accelerates the material to a velocity of
@@
A third justification for a constant shock speed as a function of radius are the models of Bertschinger and Chevalier (1985), who examined atmospheres of Mira variables supported by a periodic train of outgoing shocks. In the plane-parallel limit, they showed that periodic solutions have shock velocities that are constant as a function of height, while the density falls exponentially (when averaging over time), as shown by the calculation below. The plane-parallel limit applies when the separation between subsequent shocks is less than the stellar radius, which is only marginally true in our case. Simulations (e.g., Goldberg et al. 2022a) show that shocks are launched approximately on the convective turnover time $t_{\rm con}\sim 0.5R/v_{\rm con}$ such that the separation between shocks is $\sim 0.5R$.
@@ 2.2 Chromospheric Structure
This is similar to the structure of an isothermal atmosphere, but with the sound speed replaced by $v_{\rm s}$, and with the extra geometric factor of $(R/r)^{2}$ due to the use of the divergence of the momentum flux rather than a turbulent pressure. Just above the stellar surface, equation 5 is identical to the density profile found by Bertschinger and Chevalier (1985) for a periodic train of shocks in a plane-parallel atmosphere. Importantly, the density profile depends only on the amplitude of shocks, and not on their temporal frequency.
@@
Figure 2 shows the density profiles of equation 5 and equation 2.2 above a RSG photosphere. The underlying model is a $13.8\,M_{\odot}$ RSG with $R=844R_{\odot}$ generated using MESA (Paxton et al., 2011, 2013, 2015). This model had a ZAMS mass of $15\,M_{\odot}$ and has been evolved to core-collapse. We use MLT with $\alpha_{\rm MLT}=2.5$ to measure the convective velocity, which we evaluate where the optical depth is $\tau=P_{\rm rad}c/(Pv_{\rm con})$, where $P_{\rm rad}$ and $P$ are the radiation and total pressures. This yields approximately correct RMS convective velocities (see Goldberg et al. 2022a), which for RSGs are quite similar to the sound speed $c_{s}$. This model has $v_{\rm con}\simeq 0.08v_{\rm esc}$ near the star’s surface.
@@ 3.1 Dust Formation and Wind Acceleration
We choose to use observed wind velocities $v_{\infty}$ to constrain $x_{\rm d}$, but one could use dust opacities and mass fractions to calculate $x_{\rm d}$ and then predict the corresponding value of $v_{\infty}$. From equation 14, we have $x_{\rm d}=(R_{\rm d}/GM)(v_{\infty}^{2}-v_{\rm d}^{2})$. In this work, we use $v_{\infty}=30$ km/s, in line with observational estimates (Sjouwerman et al., 1998; van Loon et al., 2001; Marshall et al., 2004), although this velocity likely depends weakly on the stellar luminosity (e.g., Mauron and Josselin 2011). This requires $x_{\rm d}\sim 2$, approximately what we expect for realistic dust-laden gas opacities of $\kappa_{\rm d}\sim 5\,{\rm cm^{2}}/{\rm g}$ (Höfner et al., 2003).
@@ 3.2 Red Supergiant Mass Loss Rates
We implement the mass loss rate of the previous section (equation 3.1) into stellar evolution models generated with MESA as described above. At surface temperatures above $4500\,{\rm K}$ we use the “Dutch” wind prescription with scaling factor $\eta=0.5$. At cooler surface temperatures, we use our wind prescription. Our MESA inlists and run_star_extras files are available at this link.
@@
Our mass loss rates are similar for a given luminosity to those recently measured for RSGs in the LMC (Antoniadis et al., 2024; Wen et al., 2024). These mass loss rates are substantially smaller than those measured in the SMC (Yang et al., 2023), but those may be systematically overestimated according to Antoniadis et al. 2024. Most importantly, the predicted scaling of mass loss with luminosity above $L>3\times 10^{4}\,L_{\odot}$ is quite similar between our models and all of these measurments. Interestingly, the LMC/SMC measurements indicate a kink at $L\sim 3\times 10^{4}\,L_{\odot}$, with the mass loss rate flattening at smaller luminosities, which does not happen in our models. The observed mass loss rates at low luminosities are closer to the predictions of a Reimer’s mass loss prescription. It is likely that a different mass loss mechanism (e.g., mass loss driven by coronal heating due to Alfvén waves, Cranmer and Saar 2011) dominates at low luminosities.
@@
We do not plot the mass loss prescription of Schröder and Cuntz (2005), which lies between the Reimers and de Jager prescriptions for our models.
@@ 3.3 Implications of mass loss
Our massive models can exhibit exponentially increasing mass loss rate with time. As these stars lose mass, their luminosities/radii remain nearly constant but their masses decrease, decreasing $v_{\rm esc}$ and increasing the mass loss rate, creating a runaway process. Combined with the rarity of such stars due to the IMF, this implies a correspondingly small number of observed very luminous RSGs with large mass loss rates, perhaps consistent with their low observed numbers (Beasor and Smith, 2022). In contrast, the Reimers prescription predicts much lower mass loss rates, longer RSG durations, and a larger number of observed very luminous RSGs. The boil-off phenomenon of our models can thus help explain the absence of RSGs observed with luminosities larger than $L_{\rm max}\approx 10^{5.5}\,L_{\odot}$ (Davies et al., 2018), while simultaneously preserving low-mass loss rates of less massive RSGs.
Vink and Sabhahit (2023) suggested that the upward kink in mass loss rates at high luminosity ($L\gtrsim 10^{4.5}\,L_{\odot}$) is caused by a transition from an optically thin to optically thick wind. This is motivated by an analogy with line-driven winds from hot stars, which show a similar feature as photons begin to scatter multiple times in optically thick winds, allowing them to drive higher mass loss rates. However, as far as we are aware, there is no model of radiatively driven winds in cool stars, e.g., due to photon pressure on molecular absorption lines. While photon pressure on dust helps accelerate RSG winds, the mass loss rate may instead be determined by how much mass can be uplifted to the dust formation radius, so it is not clear that optically thick dust absorption would actually drive higher mass loss rates. Our model can explain the dependence of mass loss on luminosity above the kink, but it cannot explain the dependence below the kink. We speculate that a different mechanism with weaker luminosity dependence dominates mass loss at luminosities below $10^{4.5}\,L_{\odot}$. Future work will be necessary to distinguish between our suggestion and that of Vink and Sabhahit (2023).
@@ 4 Effect on Type II-P Supernovae
Our model predicts typical densities of $\rho\gtrsim 10^{-14}\,{\rm g}/{\rm cm}^{3}$ within three stellar radii of RSGs. This is a few orders of magnitude larger than expected for constant velocity winds with similar mass loss rates (Figure 2). Therefore, we expect signs of CSM interaction in type-II SNe, even for “ordinary” pre-SN mass loss rates of $\sim\!10^{-6}-10^{-5}\,M_{\odot}/{\rm yr}$. As mentioned above, the total CSM mass of our models is in the range of $10^{-3}-10^{-1}\,M_{\odot}$, slightly smaller than estimates based on light-curve modeling of type-IIP SNe exhibiting signs of CSM interaction (Morozova et al., 2017; Dessart et al., 2017; Moriya et al., 2018). Our chromospheres qualitatively resemble the less massive but more extended CSM structure needed to match both the light curve and narrow emission lines of type II-P (Dessart and Jacobson-Galán, 2023), but detailed spectral modeling will be needed to determine whether they can reproduce observations. Here we model the impact of the chromosphere on the light curves of type II-P SNe.
@@
We do not perform spectral modeling here, but we note that non-LTE modeling of flash-ionized emission lines has constrained outflow velocities to low values of $v_{\rm wind}\lesssim 30\,{\rm km/s}$ in the case of SN 2013cu (e.g., Gräfener and Vink 2016). That event was a hydrogen-poor type-IIb SN for which our models are not directly applicable. In any case, our quasi-hydrostatic chromosphere models predict low outflow/shock speeds within the dust formation radius of exploding type II-P SN progenitors, corresponding to narrow widths of $v\lesssim 30\,{\rm km/s}$ of flash-ionized emission lines, although the widths may be increased by acceleration due to radiation pressure from the subsequent SN (Tsuna et al., 2023).
@@ 4.1 Light curve modeling
When the SN shock propagates an extended dense CSM (e.g., the case for $v_{\rm s}/v_{\rm esc}=0.24$), the shock downstream is compressed into a thin dense shell due to radiative cooling. By default, SNEC obtains the “observed luminosity” by measuring the local luminosity at the photospheric radius where $\tau=2/3$. However, due to the thin-shell nature of the shocked region, we find numerical instabilities in the luminosity after the photosphere enters the shocked region (e.g., Fig 4 of Tsuna et al., 2023). To avoid this, we instead measure the bolometric luminosity at the outermost part of the computational region ($\approx 3\times 10^{16}$ cm) where photons are free-streaming ($\tau\ll 2/3$). We use this luminosity and the photospheric radius, and adopt the same bolometric corrections as done in SNEC to obtain the V-band light curves.
@@
We note that the model V-band light curves can be affected by SNEC’s assumption of gas-radiation equilibrium and a fully thermalized radiation field. We may expect departures from thermal spectra for the luminosity generated around shock breakout, and that powered by prolonged CSM interaction (e.g., Nakar and Sari, 2010; Svirski et al., 2012; Morozova et al., 2018; Haynie and Piro, 2021; Tsuna et al., 2021). On the other hand, for the shock cooling phase after breakout the thermalization is expected to be more complete (Morozova et al., 2018). Additionally, the morphology of the light curve around the peak is possibly affected by the relatively low resolution within the breakout layer ($\tau\approx$ several tens), most notably for the ones with densest CSM.